OpenStudy (anonymous):

In the right column, the measures of angles are given. match angle measure from the left column and match it with the measure of its supplement left column: 175 21 32 38 164 101 90 121 111 137 right column: 148 43 59 142 16 5 69 159 79 90 *help please*

OpenStudy (anonymous):

sum of both angles shud b 180 deg as they r supplemntry 175 and 5 ( 180-175 = 5) 21 and 159 (180 - 21 = 159) 32 and 148 (180 - 32 = 148) now u can calculate the rest on your own !!!

OpenStudy (anonymous):

@Harkirat please help me

OpenStudy (anonymous):

which question???

OpenStudy (anonymous):

thanks you!

OpenStudy (anonymous):

pls click GOOD ANSWER for me ( :})

OpenStudy (anonymous):

i did

OpenStudy (anonymous):

thanks !!

OpenStudy (anonymous):

6 & 8, there the last questions I've asked.

OpenStudy (anonymous):

ok ill check for them

OpenStudy (anonymous):

Thank you.